# Adaptive Federated Learning with Auto-Tuned Clients

### v1

@@ Abstract
Federated learning (FL) is a distributed machine learning framework where the global model of a central server is trained via multiple collaborative steps by participating clients without sharing their data. While being a flexible framework, where the distribution of local data, participation rate, and computing power of each client can greatly vary, such flexibility gives rise to many new challenges, especially in the hyperparameter tuning on both the server and the client side. We propose $\Delta$-SGD, a simple step size rule for SGD that enables each client to use its own step size by adapting to the local smoothness of the function each client is optimizing. We provide theoretical and empirical results where the benefit of the client adaptivity is shown in various FL scenarios. In particular, our proposed method achieves TOP-1 accuracy in 73% and TOP-2 accuracy in 100% of the experiments considered without additional tuning.
@@ 1 Introduction
Federated learning (FL) is a distributed machine learning framework that allows multiple clients to learn a global model collaboratively. Each client trains the model on its local data and sends only the updated model parameters to a central server for aggregation. Mathematically, FL aims to solve the following optimization problem:
@@
A key property of FL is its flexibility in how different clients participate in the overall training: the number of clients $m$, their participation rates, and computing power available to each client can all vary and change at any point during the overall training procedure. On top of that, each client’s local data is not shared with other clients or the server, resulting in certain inherent data privacy McMahan et al. (2017); Agarwal et al. (2018).
@@
While advantageous, such flexibility and data privacy also introduce a plethora of new challenges, notably: $i)$ how the server aggregates the local information coming from each client, and $ii)$ how to make sure each client meaningfully “learns” using their local data and computing device. The first challenge was partially addressed in Reddi et al. (2021), where adaptive optimization methods, such as Adagrad Duchi et al. (2011) or Adam Kingma and Ba (2014), were utilized in the aggregation step. However, the second challenge remains largely unaddressed as we elaborate below, other than the exception of Wang et al. (2021); Xie et al. (2019) reviewed in Section 2.
@@
Local data of each client is not shared, which intrinsically implies heterogeneity in terms of the size and the distribution $\mathcal{D}_{i}$ of local datasets. That is, $\mathcal{D}_{i}$ differs for each client $i$, as well as the number of samples $z\sim\mathcal{D}_{i}$. Consequently, $f_{i}(x)$ can vastly differ from client to client, making Problem (1) hard to solve (i.e., find an approximate minimizer). Moreover, the sheer amount of local updates is far larger than the number of aggregation steps, given that less communication is desired (i.e., less server-side aggregation) due to the high communication cost—typically 3-4$\times$ orders of magnitude more expensive than local computation—in distributed optimization settings Lan et al. (2020).
Figure 1: (A): CIFAR-10 classification task trained on ResNet-18. (B): MNIST classification task trained on shallow CNN. (C): CIFAR-100 classification task trained on ResNet-50. All experiments use FedAvg McMahan et al. (2017) as the server optimizer.
@@
Extensive fine-tuning of the client optimizers is often required to achieve good performance. For instance, experimental results of the famous FedAvg algorithm were obtained after performing a grid-search of typically 11-13 step sizes of the clients’ SGD (McMahan et al., 2017, Section 3), as SGD (and its variants) are highly sensitive to the step size Moulines and Bach (2011); Toulis and Airoldi (2017); Assran and Rabbat (2020); Kim et al. (2022). Similarly, in Reddi et al. (2021), 6 different client step sizes were grid-searched for different tasks, and not surprisingly, each task requires a different client step size to obtain the best result, regardless of the server-side adaptivity (Reddi et al., 2021, Table 8). Importantly, even these “fine-tunings” are done under the setting that all clients use the same step size, which is sub-optimal, given that $f_{i}$ can be vastly different per client; we analyze this further in Section 2.
@@
Initial examination as motivation. The implication of not properly tuning the client optimizer is highlighted in Figure 1. We plot the progress of test accuracies for different client optimizers, where, for all the other test cases, we intentionally use the same step size rules that were fine-tuned for the task in Figure 1(A). There, we train a ResNet-18 for CIFAR-10 dataset classification within an FL setting, where the best step sizes were used for all algorithms after grid-search; we defer the experimental details to Section 4. Hence, all methods perform reasonably well, although $\Delta$-SGD, our proposed method, achieves noticeably better test accuracy when compared to non-adaptive SGD variants –e.g., a 5% difference in final classification accuracy from SGDM– and comparable final accuracy only with adaptive SGD variants, like Adam and Adagrad.
@@
In Figure 1(B), we train a shallow CNN for MNIST data classification using the same step size rules. MNIST classification is now considered an “easy” task, and therefore SGD with the same constant and decaying step sizes from Figure 1(A) works well. However, with momentum, SGDM exhibits highly oscillating behavior, which results in slow progress and poor final accuracy, especially without decaying the step size. Adaptive optimizers, like Adam and Adagrad, show similar behavior, falling short in achieving good final accuracy, compared to their performance in the case of Figure 1(A).
@@
Similarly, in Figure 1(C), we plot the test accuracy progress for CIFAR-100 classification trained on ResNet-50, again using the same step size rules as before. Contrary to Figure 1(B), SGD with momentum (SGDM) works better than SGD, both with the constant and the decaying step sizes. Adam becomes a “good optimizer” again, but its “sibling”, Adagrad, performs worse than SGDM. On the contrary, our proposed method, $\Delta$-SGD, which we introduce in Section 3, achieves superior performance in all cases without any additional tuning.
@@
The above empirical observations beg answers to important and non-trivial questions in training FL tasks using variants of SGD methods as the client optimizer: Should the momentum be used? Should the step size be decayed? If so, when? Unfortunately, Figure 1 indicates that the answers to these questions highly vary depending on the setting; once the dataset itself or how the dataset is distributed among different clients changes, or once the model architecture changes, the client optimizers have to be properly re-tuned to ensure good performance. Perhaps surprisingly, the same holds for adaptive methods like Adagrad Duchi et al. (2011) and Adam Kingma and Ba (2014).
@@
Our hypothesis and contributions. Our paper takes a stab in this direction: we propose DELTA-SGD (DistributEd LocaliTy Adaptive SGD), a simple adaptive SGD scheme, that does not require any knowledge of the problem constants, such as the $L$-smoothness parameter. We will refer to our algorithm as $\Delta$-SGD in the rest of the text. Our adaptive step size is extended from Malitsky and Mishchenko (2020), where the focus was centralized convex optimization; we defer the readers to Section 3 for details. Our contributions can be summarized as follows:
@@
  • We extend the (centralized) adaptive step size from Malitsky and Mishchenko (2020) to the FL setting. The implication of this adaptive step size in FL is twofold: $i)$ each client can use its own step size, and $ii)$ each client’s step size adapts to the local smoothness of $f_{i}$—hence LocaliTy Adaptive— and can even increase during local iterations. Moreover, thanks to the simplicity of the proposed step size, our method is agnostic to the server optimizer as well as the loss function, and thus can be easily combined with server adaptive methods such as FedAdam Reddi et al. (2021), or methods that use different (regularized) loss functions such as FedProx Li et al. (2020) or MOON Li et al. (2021).
@@
  • We provide convergence analysis of $\Delta$-SGD in a general nonconvex setting (Theorem 1). We also provide the convergence analysis in the convex setting, extending the analysis from Malitsky and Mishchenko (2020) to a distributed one, as well as a simple centralized nonconvex convergence analysis, which was an open question posed in Malitsky and Mishchenko (2020). Due to space constraints, we only state the result of general nonconvex case in Theorem 1, and defer other theorems and proofs to Appendix A.
@@
  • We evaluate our approach on several benchmark datasets and demonstrate that $\Delta$-SGD achieves superior performance compared to other state-of-the-art FL methods. Our experiments show that our method can effectively adapt the client step size to the underlying data distribution and achieve faster convergence of the global model, without any additional tuning. Our approach can help overcome the client step size tuning challenge in FL and enable more efficient and effective collaborative learning in distributed systems.
@@ 2 Preliminaries and Related Work
A bit of background on optimization theory. Arguably the most fundamental optimization algorithm for minimizing a function $f(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the gradient descent (GD) Cauchy et al. (1847), which iterates with the step size $\eta_{t}$ as: $x_{t+1}=x_{t}-\eta_{t}\nabla f(x_{t})$. Under the assumption that $f$ is globally $L$-smooth, that is:
$\|\nabla f(x)-\nabla f(y)\|\leqslant L\cdot\|x-y\|\quad\forall x,y,$ (2)
@@
the step size $\eta_{t}=\frac{1}{L}$ is the “optimal” step size for GD, which converges at the rate, which is not improvable Nesterov (2003); Drori and Teboulle (2014), for convex $f$:
$f(x_{t+1})-f(x^{\star})\leqslant\frac{L\|x_{0}-x^{\star}\|^{2}}{2(2t+1)},$ (3)
while $\eta_{t}\in(0,\frac{2}{L})$ defines the “admissible” step size range required to ensure that GD converges for this function class Polyak (1963). Similarly, if $f$ is also (globally) $\mu$-strongly convex, that is:
$f(x)\geqslant f(y)+\langle\nabla f(y),x-y\rangle+\tfrac{\mu}{2}\|x-y\|^{2}~{}~{}\forall x,y,$
then the optimal step size is $\eta_{t}=\frac{2}{\mu+L}$ Nesterov (2003), and the rate improves to a linear one:
$\|x_{t+1}-x^{\star}\|^{2}\leqslant\Big{(}\frac{L-\mu}{L+\mu}\Big{)}^{t}\|x_{0}-x^{\star}\|^{2}.$
@@
Limitations of popular step sizes. In practice, however, the constants $L$ and $\mu$ are rarely known Boyd et al. (2004). As such, there has been a plethora of efforts in the optimization community to come up with a general step size rule or method that does not require a priori knowledge of problem constants Nesterov (2015); Orabona and Pál (2016); Levy et al. (2018), so that an optimization algorithm can work as “plug-and-play.”
@@
This is a well-known research question, with some major lines of work on this direction being: $i)$ line-search Armijo (1966); Paquette and Scheinberg (2020), $ii)$ adaptive learning rate (e.g., Adagrad Duchi et al. (2011) and Adam Kingma and Ba (2014)), and $iii)$ Polyak step size Polyak (1969); Hazan and Kakade (2019), to name a few. Unfortunately, the search for the “perfect” step size is still active Loizou et al. (2021); Defazio et al. (2022); Defazio and Mishchenko (2023); Bernstein et al. (2023); Latafat et al. (2023), as the aforementioned approaches have limitations. For instance, to use line search, solving a sub-routine is often required, inevitably incurring additional evaluations of the function and/or the gradient. To use methods like Adagrad and Adam, the knowledge of the initial distance from the solution set $D$, which is nontrivial to find, is required to ensure good performance Defazio and Mishchenko (2023); see also Section 4. Lastly, Polyak step size requires a sharp knowledge of $f(x^{\star})$ Hazan and Kakade (2019), which is not always known a priori, and is unclear what value to use for an arbitrary loss function and model.
@@
Implications in FL scenarios. On top of that, even with the knowledge of $L$, GD can be slow. Indeed, the global $L$-smoothness condition, as described above, needs to be satisfied for all $x$ and $y$, implying even finite $L$ can be arbitrarily large. This results in a small step size, leading to slow convergence, as seen in (3).
@@
Naturally, the challenge is aggravated in the distributed case. To solve (1), under the assumption that each $f_{i}$ is $L_{i}$-smooth, the step size of the form ${1}/{L_{\text{max}}}$ where ${L_{\text{max}}}:=\max_{i}L_{i}$ is often used Yuan et al. (2016); Scaman et al. (2017); Qu and Li (2019), to ensure convergence Uribe et al. (2020). Yet, one can easily imagine a situation where $L_{i}\gg L_{j}$ for $i\neq j,$ in which case the convergence of $f_{j}(x)$ can be arbitrarily slow by using step size ${1}/{L_{\text{max}}}$.
There are plenty of practical examples where such a scenario could happen. For instance, imagine smartphone photo albums of the members of a fishing club versus a bird-watching club. The albums of these two groups could have vastly different data distributions, one skewed with pictures including fishes and rivers, while the other skewed with birds and forests Wah et al. (2011). On top of that, smartphones from different manufacturers nowadays have distinctly different camera filters applied—recently a big selling point for companies like Apple, introducing technologies like “portrait mode,” and Samsung, introducing 100$\times$ zoom capability, both based on image processing techniques—making even similar pictures to have distinct pixel values, when taken with different smartphones.
@@
Therefore, to ensure that each client learns useful information in the FL setting as in (1), ideally: $i)$ each agent should be able to use its own step size, instead of crude ones like $1/L_{\text{max}}$ for all; $ii)$ the individual step size should be “locally adaptive” to the function $f_{i},$ (i.e., even $1/L_{i}$ can be too crude, analogously to the centralized GD); and $iii)$ the step size should not depend on problem constants like $L_{i}$ an $\mu_{i}$, which are often unknown. In Section 3, we introduce DistributEd LocaliTy Adaptive SGD ($\Delta$-SGD), which satisfies all of the aforementioned desiderata.
Related work on FL. FedAvg McMahan et al. (2017) is the simplest and probably the most famous FL algorithm, which takes a (weighted) average of the client parameters in the aggregation step after clients perform some local iterations. While FedAvg has achieved great results empirically, it has been observed recently that it can suffer from model drift Praneeth Karimireddy et al. (2019) induced by the heterogeneity Hsu et al. (2019). At the core, simple averaging is too crude of an operation. As such, FedAvg uses a weighted average heuristic, where the weights are based on the number of data $n_{i}$ for each client $i$ McMahan et al. (2017).
@@
The authors in Reddi et al. (2021) took this idea one step further: it showed that FedAvg can be seen as a special case where both the clients and the server use the SGD optimizer, with a server learning rate of 1. To handle the heterogeneity and the model drift, they proposed to use adaptive optimization scheme at the server, resulting in, e.g., FedAdam or FedYogi, which respectively use Adam Kingma and Ba (2014) and Yogi Zaheer et al. (2018) as the server optimizer, and can improve upon FedAvg in various scenarios. Note that this fundamentally differs from our approach, where we propose a client adaptive optimizer that can be easily combined with sophisticated server-side aggregation methods like FedAdam.
Perhaps the closest works in spirit to our work are AdaAlter Xie et al. (2019) and locality adaptivity in FL Wang et al. (2021). Both works utilize an adaptive client optimizer, similar to Adagrad Duchi et al. (2011), with either delayed or restarted updates on the adaptive variables (e.g., pre-conditioners and momentum buffers). However, as noted in Reddi et al. (2021), AdaAlter incurs twice the memory compared to FedAvg, as AdaAlter requires communicating both the model parameters as well as the “client accumulators,” which is of the same dimension as model parameters. Wang et al. (2021) overcomes this issue by simply restarting the client optimizer states, but requires complicated local and global correction step to achieve good performance. On the contrary, $\Delta$-SGD changes virtually nothing compared to FedAvg other than the simple step size computation, and yet still achieves impressive performance; see Section 4.
@@
Other approaches have handled the heterogeneity by changing the loss function. FedProx Li et al. (2020) adds an $\ell_{2}$-norm proximal term to the individual loss functions such that the local parameters do not diverge too much from the previously communicated global model. Similarly, MOON Li et al. (2021) uses the cross-entropy loss between the current and previous models. As mentioned in the introduction, our proposed method can be seamlessly combined with these approaches.
@@
Another line of works attempts to handle heterogeneity by utilizing control-variates Praneeth Karimireddy et al. (2019); Liang et al. (2019); Karimireddy et al. (2020), a similar idea to variance reduction from single-objective optimization Johnson and Zhang (2013); Gower et al. (2020). While theoretically appealing, these methods either require full gradient computation periodically or are not usable when not all the clients participate, as the clients must maintain a state throughout all rounds.

### v3

@@ Abstract
Federated learning (FL) is a distributed machine learning framework where the global model of a central server is trained via multiple collaborative steps by participating clients without sharing their data. While being a flexible framework, where the distribution of local data, participation rate, and computing power of each client can greatly vary, such flexibility gives rise to many new challenges, especially in the hyperparameter tuning on the client side. We propose $\Delta$-SGD, a simple step size rule for SGD that enables each client to use its own step size by adapting to the local smoothness of the function each client is optimizing. We provide theoretical and empirical results where the benefit of the client adaptivity is shown in various FL scenarios.
@@ 1 Introduction
Federated Learning (FL) is a machine learning framework that enables multiple clients to collaboratively learn a global model in a distributed manner. Each client trains the model on their local data, and then sends only the updated model parameters to a central server for aggregation. Mathematically, FL aims to solve the following optimization problem:
@@
A key property of FL its flexibility in how various clients participate in the overall training procedure. The number of clients, their participation rates, and computing power available to each client can vary and change at any time during the training. Additionally, the local data of each client is not shared with others, resulting in better data privacy (McMahan et al., 2017; Agarwal et al., 2018).
@@
While advantageous, such flexibility also introduces a plethora of new challenges, notably: $i)$ how the server aggregates the local information coming from each client, and $ii)$ how to make sure each client meaningfully “learns” using their local data and computing device. The first challenge was partially addressed in Reddi et al. (2021), where adaptive optimization methods such as Adam (Kingma & Ba, 2014) was utilized in the aggregation step. Yet, the second challenge remains largely unaddressed.
@@
Local data of each client is not shared, which intrinsically introduces heterogeneity in terms of the size and the distribution of local datasets. That is, $\mathcal{D}_{i}$ differs for each client $i$, as well as the number of samples $z\sim\mathcal{D}_{i}$. Consequently, $f_{i}(x)$ can vastly differ from client to client, making the problem in (1) hard to optimize. Moreover, the sheer amount of local updates is far larger than the number of aggregation steps, due to the high communication cost—typically 3-4$\times$ orders of magnitude more expensive than local computation—in distributed settings (Lan et al., 2020).
Figure 1: Illustration of the effect of not properly tuning the client step sizes. In (A), each client optimizer uses the best step size from grid-search. Then, the same step size from (A) is intentionally used in settings (B) and (C). Only $\Delta$-SGD works well across all settings without additional tuning.
@@
As a result, extensive fine-tuning of the client optimizers is often required to achieve good performance in FL scenarios. For instance, experimental results of the well-known FedAvg algorithm were obtained after performing a grid-search of typically 11-13 step sizes of the clients’ SGD (McMahan et al., 2017, Section 3), as SGD (and its variants) are highly sensitive to the step size (Toulis & Airoldi, 2017; Assran & Rabbat, 2020; Kim et al., 2022b). Similarly, in Reddi et al. (2021), 6 different client step sizes were grid-searched for different tasks, and not surprisingly, each task requires a different client step size to obtain the best result, regardless of the server-side adaptivity (Reddi et al., 2021, Table 8). Importantly, even these “fine-tunings” are done under the setting that all clients use the same step size, which is sub-optimal, given that $f_{i}$ can be vastly different per client; we analyze this further in Section 2.
@@
Initial examination as motivation. The implication of not properly tuning the client optimizer is highlighted in Figure 1. We plot the progress of test accuracies for different client optimizers, where, for all the other test cases, we intentionally use the same step size rules that were fine-tuned for the task in Figure 1(A). There, we train a ResNet-18 for CIFAR-10 dataset classification within an FL setting, where the best step sizes for each (client) optimizer was used after grid-search; we defer the experimental details to Section 4. Hence, all methods perform reasonably well, although $\Delta$-SGD, our proposed method, achieves noticeably better test accuracy when compared to non-adaptive SGD variants –e.g., a 5% difference in final classification accuracy from SGDM– and comparable final accuracy only with adaptive SGD variants, like Adam and Adagrad.
@@
In Figure 1(B), we train a shallow CNN for MNIST classification, using the same step sizes from (A). MNIST classification is now considered an “easy” task, and therefore SGD with the same constant and decaying step sizes from Figure 1(A) works well. However, with momentum, SGDM exhibits highly oscillating behavior, which results in slow progress and poor final accuracy, especially without decaying the step size. Adaptive optimizers, like Adam and Adagrad, show similar behavior, falling short in achieving good final accuracy, compared to their performance in the case of Figure 1(A).
@@
Similarly, in Figure 1(C), we plot the test accuracy progress for CIFAR-100 classification trained on a ResNet-50, again using the same step size rules as before. Contrary to Figure 1(B), SGD with momentum (SGDM) works better than SGD, both with the constant and the decaying step sizes. Adam becomes a “good optimizer” again, but its “sibling,” Adagrad, performs worse than SGDM. On the contrary, our proposed method, $\Delta$-SGD, which we introduce in Section 3, achieves superior performance in all cases without any additional tuning.
@@
The above empirical observations beg answers to important and non-trivial questions in training FL tasks using variants of SGD methods as the client optimizer: Should the momentum be used? Should the step size be decayed? If so, when? Unfortunately, Figure 1 indicates that the answers to these questions highly vary depending on the setting; once the dataset itself or how the dataset is distributed among different clients changes, or once the model architecture changes, the client optimizers have to be properly re-tuned to ensure good performance. Perhaps surprisingly, the same holds for adaptive methods like Adagrad (Duchi et al., 2011) and Adam (Kingma & Ba, 2014).
@@
Our hypothesis and contributions. Our paper takes a stab in this direction: we propose DELTA-SGD (DistributEd LocaliTy Adaptive SGD), a simple adaptive distributed SGD scheme, that can automatically tune its step size based on the available local data. We will refer to our algorithm as $\Delta$-SGD in the rest of the text. Our contributions can be summarized as follows:
@@
  • We propose $\Delta$-SGD, which has two implications in FL settings: $i)$ each client can use its own step size, and $ii)$ each client’s step size adapts to the local smoothness of $f_{i}$ –hence LocaliTy Adaptive– and can even increase during local iterations. Moreover, due to the simplicity of the proposed step size, $\Delta$-SGD is agnostic to the loss function and the server optimizer, and thus can be combined with methods that use different loss functions such as FedProx (Li et al., 2020) or MOON (Li et al., 2021), or adaptive server methods such as FedAdam (Reddi et al., 2021).
@@
  • We provide convergence analysis of $\Delta$-SGD in a general nonconvex setting (Theorem 1). We also prove convergence in convex setting; due to space constraints, we only state the result of the general nonconvex case in Theorem 1, and defer other theorems and proofs to Appendix A.
@@
  • We evaluate our approach on several benchmark datasets and demonstrate that $\Delta$-SGD achieves superior performance compared to other state-of-the-art FL methods. Our experiments show that $\Delta$-SGD can effectively adapt the client step size to the underlying local data distribution, and achieve convergence of the global model without any additional tuning. Our approach can help overcome the client step size tuning challenge in FL and enable more efficient and effective collaborative learning in distributed systems.
@@ 2 Preliminaries and Related Work
A bit of background on optimization theory. Arguably the most fundamental optimization algorithm for minimizing a function $f(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the gradient descent (GD), which iterates with the step size $\eta_{t}$ as: $x_{t+1}=x_{t}-\eta_{t}\nabla f(x_{t})$. Under the assumption that $f$ is globally $L$-smooth, that is:
$\|\nabla f(x)-\nabla f(y)\|\leqslant L\cdot\|x-y\|\quad\forall x,y\in\mathbb{R}^{d},$ (2)
@@
the step size $\eta_{t}=\frac{1}{L}$ is the “optimal” step size for GD, which converges for convex $f$ at the rate:
$f(x_{t+1})-f(x^{\star})\leqslant\frac{L\|x_{0}-x^{\star}\|^{2}}{2(2t+1)}.$ (3)
@@
Limitations of popular step sizes. In practice, however, the constants like $L$ are rarely known (Boyd et al., 2004). As such, there has been a plethora of efforts in the optimization community to develop a universal step size rule or method that does not require a priori knowledge of problem constants so that an optimization algorithm can work as “plug-and-play” (Nesterov, 2015; Orabona & Pál, 2016; Levy et al., 2018). A major lines of work on this direction include: $i)$ line-search (Armijo, 1966; Paquette & Scheinberg, 2020), $ii)$ adaptive learning rate (e.g., Adagrad (Duchi et al., 2011) and Adam (Kingma & Ba, 2014)), and $iii)$ Polyak step size (Polyak, 1969), to name a few.
@@
Unfortunately, the pursuit of finding the “ideal” step size is still active (Loizou et al., 2021; Defazio & Mishchenko, 2023; Bernstein et al., 2023), as the aforementioned approaches have limitations. For instance, to use line search, solving a sub-routine is required, inevitably incurring additional evaluations of the function and/or the gradient. To use methods like Adagrad and Adam, the knowledge of $D$, the distance from the initial point to the solution set, is required to ensure good performance (Defazio & Mishchenko, 2023). Similarly, Polyak step size requires knowledge of $f(x^{\star})$ (Hazan & Kakade, 2019), which is not always known a priori, and is unclear what value to use for an arbitrary loss function and model; see also Section 4 where not properly estimating $D$ or $f(x^{\star})$ resulting in suboptimal performance of relevant client optimizers in FL scenarios.
@@
Implications in distributed/FL scenarios. The global $L$-smoothness in (2) needs to be satisfied for all $x$ and $y$, implying even finite $L$ can be arbitrarily large. This results in a small step size, leading to slow convergence, as seen in (3). Malitsky & Mishchenko (2020) attempted to resolve this issue by proposing a step size for (centralized) GD that depends on the local smoothness of $f$, which by definition is smaller than the global smoothness constant $L$ in (2). $\Delta$-SGD is inspired from this work.
@@
Naturally, the challenge is aggravated in the distributed case. To solve (1) under the assumption that each $f_{i}$ is $L_{i}$-smooth, the step size of the form ${1}/{L_{\text{max}}}$ where ${L_{\text{max}}}:=\max_{i}L_{i}$ is often used (Yuan et al., 2016; Scaman et al., 2017; Qu & Li, 2019), to ensure convergence (Uribe et al., 2020). Yet, one can easily imagine a situation where $L_{i}\gg L_{j}$ for $i\neq j,$ in which case the convergence of $f_{j}(x)$ can be arbitrarily slow by using step size of the form ${1}/{L_{\text{max}}}$.
@@
Therefore, to ensure that each client learns useful information in FL settings as in (1), ideally: $i)$ each agent should be able to use its own step size, instead of crude ones like $1/L_{\text{max}}$ for all agents; $ii)$ the individual step size should be “locally adaptive” to the function $f_{i},$ (i.e., even $1/L_{i}$ can be too crude, analogously to the centralized GD); and $iii)$ the step size should not depend on problem constants like $L_{i}$ or $\mu_{i}$, which are often unknown. In Section 3, we introduce DistributEd LocaliTy Adaptive SGD ($\Delta$-SGD), which satisfies all of the aforementioned desiderata.
@@
Related work on FL. The FL literature is vast; here, we focus on the results that closely relate to our work. The FedAvg algorithm (McMahan et al., 2017) is one of the simplest FL algorithms, which averages client parameters after some local iterations. Reddi et al. (2021) showed that FedAvg is a special case of a meta-algorithm, where both the clients and the server use SGD optimizer, with the server learning rate being 1. To handle data heterogeneity and model drift, they proposed using adaptive optimizers at the server. This results in algorithms such as  FedAdam and FedYogi, which respectively use the Adam (Kingma & Ba, 2014) and Yogi (Zaheer et al., 2018) as server optimizers. Our approach is orthogonal and complimentary to this, as we propose a client adaptive optimizer that can be easily combined with these server-side aggregation methods (c.f., Appendix B.4).
@@
Other approaches have handled the heterogeneity by changing the loss function. FedProx (Li et al., 2020) adds an $\ell_{2}$-norm proximal term to to handle data heterogeneity. Similarly, MOON (Li et al., 2021) uses the model-contrastive loss between the current and previous models. Again, our proposed method can be seamlessly combined with these approaches (c.f., Table 4(b), Appendix B.6 and B.5).
The closest related work to ours is a concurrent work by Mukherjee et al. (2023). There, authors utilize the Stochastic Polyak Stepsize (Loizou et al., 2021) in FL settings. We do include SPS results in Section 4. There are some previous works that considered client adaptivity, including AdaAlter (Xie et al., 2019a) and Wang et al. (2021). Both works utilize an adaptive client optimizer similar to Adagrad. However, AdaAlter incurs twice the memory compared to FedAvg, as AdaAlter requires communicating both the model parameters as well as the “client accumulators”; similarly, Wang et al. (2021) requires complicated local and global correction steps to achieve good performance. In short, previous works that utilize client adaptivity require modification in server-side aggregation; without such heuristics, Adagrad (Duchi et al., 2011) exhibits suboptimal performance, as seen in Table 1.
@@
Lastly, another line of works attempts to handle heterogeneity by utilizing control-variates (Praneeth Karimireddy et al., 2019; Liang et al., 2019; Karimireddy et al., 2020), a similar idea to variance reduction from single-objective optimization (Johnson & Zhang, 2013; Gower et al., 2020). While theoretically appealing, these methods either require periodic full gradient computation or are not usable under partial client participation, as all the clients must maintain a state throughout all rounds.